OpenStudy (anonymous):

solve. cos(pi/4-x)cos(pi/4-y)-sin(pi/4-x)sin(pi/4-y) = sin(x+y)

OpenStudy (gorv):

cosAcosB-sinAsinB=cos(A-B) so cos(pi/4-x-(pi/4-y))=sin(x+y) =cos(pi/4-x-pi/4+y)=sin(x+y) cos(-x+y)=sin(x+y) cos(x-y)=sin(x+y) cosxcosy-sinxsiny=sinxcosy-cosxsiny cosxcosy-sinxcosy=sinxsiny-cosxsiny cosy(cosx-sinx)=siny(sinx-cosx) cosy(cosx-sinx)-siny(sinx-cosx)=0 cosy(cosx-sinx)+siny(cosx-sinx)=0 (cosx-sinx)(cosy+siny)=0 cosx-sinx=0 or cosy+siny=0